OpenStudy (doodleduck8):

Which property is (1/5) [(5 + n) + -n] 1. associative 2. reflexive 3. commutative 4. distributive

zepdrix (zepdrix):

No property is shown. We need to see some type of equality taken place between two expressions in order to show a property. Associative: (1/5) [(5 + n) + -n] = (1/5) [5 + (n + -n)] Distributive: (1/5) [(5 + n) + -n] = (1/5)(5+n) + (1/5)(-n) Commutative: (1/5) [(5 + n) + -n] = [(5 + n) + -n] (1/5) Need more information in order to be able to answer the question.